### v3

Abstract
In this work, we consider the non-invasive medical imaging modality of Electrical Impedance Tomography, where the problem is to recover the conductivity in a medium from a set of data that arises out of a current-to-voltage map (Neumann-to-Dirichlet operator) defined on the boundary of the medium. We formulate this inverse problem as an operator-learning problem where the goal is to learn the implicitly defined operator-to-function map between the space of Neumann-to-Dirichlet operators to the space of admissible conductivities. Subsequently, we use an operator-learning architecture, popularly called DeepONets, to learn this operator-to-function map. Thus far, most of the operator learning architectures have been implemented to learn operators between function spaces. In this work, we generalize the earlier works and use a DeepONet to actually learn an operator-to-function map. We provide a Universal Approximation Theorem type result which guarantees that this implicitly defined operator-to-function map between the the space of Neumann-to-Dirichlet operator to the space of conductivity function can be approximated to an arbitrary degree using such a DeepONet. Furthermore, we provide a computational implementation of our proposed approach and compare it against a standard baseline. We show that the proposed approach achieves good reconstructions and outperforms the baseline method in our experiments.
keywords:
Electrical Impedance Tomography, DeepONets, Calderon’s Problem
{MSCcodes}
44A12, 65R10, 94A12
1 Introduction
In recent years tremendous progress has been made in developing accurate, non-invasive medical imaging techniques for disease detection and diagnosis and to make it suitable for practical deployment. One such medical imaging technology is Electrical Impedance Tomography (EIT). EIT is a functional imaging technique that seeks to reconstruct the spatial conductivity within a domain by applying currents and measuring voltages across electrodes attached to the boundary of the domain of interest. EIT has been shown to be clinically useful in monitoring lung function and breast cancer, see e.g. [14, 20]. It is now becoming an increasingly promising neuro-imaging method, e.g. in early detection of cerebral edema and monitoring the effect of mannitol dehydration, [74]. It is also being used in epilepsy imaging to detect seizure onset zones, [73]. Moreover, a bedside EIT stroke monitoring system has been developed for clinical practice and there is growing consensus among practitioners for utilizing EIT in multi-modal imaging for diagnosing and classifying strokes, see e.g.[2, 63]. In spite of the fact that EIT has lower spatial resolution than some other tomographic techniques such as computed tomography and MRI, it has relatively fast data acquisition times, relatively inexpensive hardware and does not use ionizing radiations, making it a versatile and safer alternative to many other medical imaging modalities, see [37] and the references therein.
Recall that a typical EIT experiment consists of injecting a weak excitation current through boundary electrodes surrounding the region of interest and imaging the distribution of electrical parameters in the region of interest by measuring voltage signals on such electrodes. Thus the measured data arises as a current-to-voltage map, which is physically related to the internal conductivity in the region of interest. Different kinds of tissues can be differentiated based on their different conductivity values, which may also be altered in diseased states. As such, the inverse problem consists of detecting physiological and pathological changes in the conductivity profile based on measurements of the current-to-voltage map across the boundary electrodes. More recently, there is a growing interest among theoreticians and practitioners alike in exploiting various machine-learning based algorithms and neural-network architectures to solve the EIT inverse problem. In this work, we use a DeepONet architecture and provide approximation-theoretic guarantees which show that by increasing the number of electrodes and by appropriately choosing the number of network parameters, one can estimate the internal conductivity to any arbitrary precision in a suitable norm. Furthermore, we implement such an architecture based on the theoretical DeepONet framework and compare its performance against a standard baseline to show that the propose neural-network outperforms that state-of-the-art baseline in our numerical experiments.
For the benefit of the readers, in the following subsections we will first give a self-contained mathematical description of the inverse problem at hand and survey some related literature. In the subsequent sections, we will describe our main theoretical results, give details about the proposed architecture and support our findings with numerical experiments.
1.1 Mathematical preliminaries and background of the EIT problem
[…]
(1) $\nabla\cdot(\gamma\nabla u)$ $=0\quad\mbox{ in }\Omega;$
(2) $\begin{split}\gamma\partial_{\nu}u|_{\partial\Omega}&=g\mbox{ on }\partial\Omega\quad(\text{or more generally }\Sigma)\\ &=0\mbox{ else.}\end{split}$
Here $\gamma\in L^{\infty}_{+}(\Omega)$ is the internal conductivity in the medium $\Omega$ such that $\gamma=1\mbox{ in }\Omega\setminus\Omega^{\prime}$, and $u(x)$ is the electric potential in the interior. Loosely speaking, we will refer to $\gamma\equiv 1$ as the background conductivity. The Neumann boundary-data models the applied current. In the abstract continuum model (CM), the corresponding boundary measurements of the voltage (Dirichlet data) are given by the Neumann-to-Dirichlet (N-t-D) operator defined on the boundary $\partial\Omega$ ( or more generally on a subset $\Sigma\subset\partial\Omega$):
(3) $\Lambda_{\gamma}:L^{2}_{\diamond}(\partial\Omega)\to L^{2}_{\diamond}(\partial\Omega),\quad g\mapsto u|_{\partial\Omega}.$
To emphasize, $u|_{\partial\Omega}$ refer to the voltage measurements on the the boundary $\partial\Omega$. Additionally, here and below we will use $\Lambda_{1}$ to denote the background N-t-D operator corresponding to the background conductivity. We would like to mention here, that in the abstract mathematical formulation, typically the Dirichlet-to-Neumann (D-t-N) map is considered instead of the N-t-D map as is considered in this work. However, in practical settings such as in EIT, the data that is generated is in the form of current-to-voltage maps which is best captured by the N-t-D maps. A more practical formulation of the EIT problem is given by the Complete Electrode Model (CEM), [18]. As before we let $\gamma\in L^{\infty}_{+}(\Omega)$. Let there be $M$ open, connected and mutually disjoint electrodes, $E_{m}\subset\partial\Omega$, $m=\{1,\cdots,M\}$ having the same contact impedance $z>0$. Assuming that the injected currents $I_{m}$ on the electrodes satisfy the Kirchoff condition, i.e. $\Sigma_{m=1}^{M}I_{m}=0$, the resulting electric potential $(u,U)\in H^{1}(\Omega)\times\mathbb{R}^{M}$ satisfies (1) along with the following conditions on the boundary:
(4) $\gamma\partial_{\nu}u$ $=0;\quad\text{on }\partial\Omega\setminus\bigcup E_{m}$
(5) $\int_{E_{m}}\gamma\partial_{\nu}udS$ $=I_{m}\quad\text{on }E_{m},m=1,\cdots,M$
(6) $u+z\sigma\partial_{\nu}u$ $=U_{m}\quad\text{on }E_{m},m=\{1,\cdots,M\}$
where the vector $U=(U_{1},\cdots,U_{M})$ represents voltage measurement on the electrodes satisfying a ground condition $\Sigma_{m=1}^{M}U_{m}=0$. We can thus define the $M-$electrode current to voltage operator, $R_{M}(\gamma):\mathbb{R}^{M}_{\diamond}\to\mathbb{R}^{M}_{\diamond}:I:=(I_{1},\cdots,I_{M})\mapsto(U_{1},\cdots,U_{M})=U$ and $U=R_{M}(\gamma)I$. For later use, we also define the ‘shifted’ CEM operator, $\tilde{R}_{M}(\gamma):={R}_{M}(\gamma)-{R}_{M}(1)$ where ${R}_{M}(1)$ is the background CEM current-to-voltage operator. Since the background operator $\tilde{R}_{M}(1)$ is known, $\tilde{R}_{M}(\gamma)$ is completely determined by $R_{M}(\gamma)$ and vice-versa. Before we define the inverse problem at hand, we will first define the set of admissible conductivities considered in this work. We recall the following definition from [28, Definition 2.2].
Definition 1.1.
A set $\mathcal{F}\subset L^{\infty}{(\Omega)}\cap L^{2}(\Omega)$ is called a finite dimensional subset of piecewise analytic functions if its linear span
$\mathrm{span}\ \mathcal{F}=\bigg{\{}\sum\limits_{j=1}^{k}\lambda_{j}f_{j}:k\in\mathbb{N},\lambda_{j}\in\mathbb{R},f_{j}\in\mathcal{F}\bigg{\}}\subset L^{\infty}(\Omega)\cap L^{2}(\Omega)$
contains only piecewise analytic function and $\mathrm{dim}\ (\mathrm{span}\ \mathcal{F})<\infty.$
Given such a finite dimensional subset $\mathcal{F}$ and two positive numbers $m,M$ we denote the space of admissible conductivities $\Gamma$ as:
(7) $\Gamma$ =$=\{\gamma\in\mathcal{F}:M>\gamma(x)\geq m>0,\gamma(x)=1\text{ in }\Omega\setminus\Omega^{\prime}\}$
As $\Omega$ is bounded, any such admissible conductivity is clearly contained in $L^{2}(\Omega).$ Furthermore, $\Gamma$ is closed by virtue of being a known finite dimensional subspace of $L^{\infty}(\Omega)\cap L^{2}(\Omega)$. Such conductivities have been considered in the literature before, see e.g. [28, 22] The inverse problem in CM can then be stated as, given all possible pairs of associated functions $(g,u|_{\Sigma})$, we want to find $\gamma\in\Gamma$.
1.2 Related Research
An EIT experiment applies the electrical current (Neumann data) on $\partial\Omega$ to measure the electrical potential differences on $\partial\Omega$. By doing so, the so-called Neumann-to-Dirichlet (NtD) operator $\Lambda_{\sigma}$ is obtained. The inverse problem is reconstructing the unknown conductivity $\sigma$ from a set of EIT experiments [7, 13, 26, 49]. Early important results by Sylvester and Uhlmann [71], and by Nachmann [56] established that unique recovery of the conductivity from the boundary data in the form of D-t-N map is possible in the (abstract) Calderón problem set-up. Stability estimates for the Calderón problem were given in [4, 58]. Optimality of such stability estimates, at least up to the order of the exponents, was proved in [52]. [66, 72] and many references therein contain an exhaustive survey of results in this area.
EIT is a highly nonlinear and strongly ill-posed problem. Therefore, in the classical literature, EIT requires a regularization for a good reconstruction [33, 34]. In the literature, there are abundant approaches for solving the EIT inverse problem including the factorization method [39], d-bar method [32], or variational methods for least-square fitting [33, 34, 3, 38] as well as, on statistical methods [1, 3, 35, 69].
There is also some recent research on deep learning for solving the EIT inverse problem. They include using Physics inspired neural networks [62], different kinds of convolutional neural networks [25, 19, 24], and shape reconstruction [70]. On the other hand, an architecture called Deep-Operator-Networks (DeepONet, or DON) was proposed in [51] for learning implicitly defined operator maps between infinite-dimensional function spaces. This architecture was a massive generalization of an earlier known (shallow) operator-network architecture proposed in [12], see also [10, 11]. Subsequently, variations of DeepONet (more generally, neural-operators) have been used for solving important physical problems and substantial efforts have been made to understand the theoretical underpinnings of why this approach seems to be particularly efficient, see e.g. [76, 29, 67, 61, 8, 36, 47, 64, 23, 6, 48, 43, 16, 41, 46, 40] and the many references therein. The main idea in operator-learning is that it seeks to directly approximate the operator map between infinite-dimensional spaces and in many cases can be shown to approximate any continuous operator to desired accuracy. Hereby, we do not have the disadvantages of the somewhat limited resolution of traditional convolutional neural-network based methods such as [25, 19, 24]. While most of the work in the literature has so far appeared in approximating a map between two function spaces, in EIT, we are interesting in learning the map between a space of (N-t-D) operators to a space of (conductivity) functions. Furthermore, we do not only provide strong experimental results for the DeepONet implementation for EIT but also provide approximation-theoretic guarantees as a theoretical justification for this approach.
In the computer-vision community, a similar DeepONet based architecture was introduced as implicit-neural-networks or, INNs. The idea here is to replace conventional discretized signal representation with implicit representations of 3D objects via neural networks. In discrete representations images are represented as discrete grids of pixels, audio signals are represented as discrete samples of amplitudes, and 3D shapes are usually parameterized as grids of voxels, point clouds, or meshes. In contrast, Implicit Neural Representations seek to represent a signal as a continuous function that maps the domain of the signal (i.e., a coordinate, such as a pixel coordinate for an image) to the value of the continuous signal at that coordinate (for an image, an R,G,B color). Some early work in this field includes single-image 3D reconstructions [53, 30, 68], representing texture on 3D objects [59], representing surface light fields in 3D [60], or 3D reconstructions from many images [54]. Both DeepONets and INNs seek to learn implicitly defined continuous maps between the input and output spaces and thus share many similarities. Motivated by these works, one of the authors of this article used such network architectures for shape reconstruction in the EIT problem [70].
2 The Learning Problem for EIT
Our goal in this article is to propose a DeepONet based neural-network architecture for learning to invert Neumann-to-Dirichlet operator relevant to the problem of EIT. To formally set up the learning problem, let us consider the shifted abstract N-t-D operator $\tilde{\Lambda}_{\gamma}={\Lambda}_{\gamma}-\Lambda_{1}$, where $\Lambda_{1}$ is the background N-t-D operator. Analogous to the CEM case, $\tilde{\Lambda}_{\gamma}$ is completely determined by $\Lambda_{\gamma}$ and vice-versa. Now we list some mapping properties of the operator $\tilde{\Lambda}_{\gamma}$ which will be crucial in our analysis. First of all, note that $\tilde{\Lambda}_{\gamma}$ is a smoothing operator, see e.g. [31, Theorem A.3] and [21, (2.5)]. In particular the operator norm of $\tilde{\Lambda}_{\gamma}$ is bounded, i.e., $\lvert\lvert\tilde{\Lambda}_{\gamma}\rvert\rvert_{\mathcal{L}(H^{s}\to H^{t})}\leq C$ for all $s,t\in\mathbb{R}$. It can be shown that the operator $\tilde{\Lambda}_{\gamma}$ belongs to the class of Hilbert-Schmidt operators. For readers’ convenience we recall the definition of a Hilbert-Schmidt operator from [1] here. Consider an orthonormal basis of $L^{2}(\partial\Omega)$ given by $\{\phi_{k}:=\phi_{k}^{(0)}\}_{k=0}^{\infty}$ consisting of real-valued eigenfunctions of the Laplace–Beltrami operator on the compact manifold $\partial\Omega$. By removing the constant function $\phi_{0}$ from the above collection, we obtain bases for $L^{2}(\partial\Omega)/\mathbb{C}\quad\text{and}\quad L^{2}_{\diamond}(\partial\Omega)=H^{0}_{\diamond}(\partial\Omega).$ Now, by an appropriately rescaling these basis functions we can also get orthonormal bases $\{\phi_{k}^{(r)}\}_{k=1}^{\infty}$ of all $H^{r}(\partial\Omega)/\mathbb{C}$ and $H^{r}_{\diamond}(\partial\Omega)$ spaces, $r\in\mathbb{R}$. For $j,k\in\mathbb{N}$, denote the tensor product operator, $b_{jk}^{(r)}:H^{r}(\partial\Omega)\to L^{2}(\partial\Omega)$
$b_{jk}^{(r)}(f)=\phi_{j}^{(r)}\otimes\phi_{k}^{(0)}(f):=\langle f,\phi_{j}^{(r)}\rangle_{H^{r}(\partial D)}\phi_{k}^{(0)},\quad f\in H^{r}(\partial\Omega),$
and define the space of linear operators
$\mathbb{H}_{r}:=\left\{T:H^{r}(\partial\Omega)\to L^{2}(\partial\Omega),\quad T=\sum_{j,k=1}^{\infty}t_{jk}b_{jk}^{(r)}:t_{jk}\in\mathbb{R},\sum_{j,k=1}^{\infty}t_{jk}^{2}<\infty\right\}.$
The elements of $\mathbb{H}_{r}$ are the ‘Hilbert–Schmidt’ operators between $H^{r}(\partial\Omega)$ and $L^{2}(\partial\Omega)$. Furthermore, $\mathbb{H}_{r}$ is itself a Hilbert space with inner product
$\langle S,T\rangle_{\mathbb{H}_{r}}:=\sum_{j,k=1}^{\infty}s_{jk}t_{jk}\equiv\sum_{j,k=1}^{\infty}\langle S\phi_{j}^{(r)},\phi_{k}^{(0)}\rangle_{L^{2}(\partial D)}\langle T\phi_{j}^{(r)},\phi_{k}^{(0)}\rangle_{L^{2}(\partial D)}.$
From [1, Lemma 18], we know that $\tilde{\Lambda}_{\gamma}\in\mathbb{H}_{r}$ for any $r$. Consider the set $D_{\Lambda}=\{\tilde{\Lambda}_{\gamma}:\gamma\in\Gamma\}$ of N-t-D operators corresponding to admissible conductivities. We show in Lemma 7.1 that $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$.
Formally, the problem is to ‘learn’ the map $G^{\text{true}}:D_{\Lambda}\to\Gamma$. More particularly, we want to design a deep neural network (DeepONet) to approximate the map, $G^{\text{true}}$. For learning operators between infinite dimensional spaces, DeepONets were introduced in [44, 51]. These generalized the (shallow-net) architecture proposed in [11, 12]. Theoretical guarantees for DeepONets were given in [44] where the neural network tries to learn operators between two Hilbert spaces. This theory was further generalized in [40, 42] and an architecture called Fourier Neural Operators was introduced for learning operators between Banach spaces having a certain approximation property. Fairly recently, DeepONets and FNOs have been synthesized as Neural Inverse Operators and they have been used effectively in solving several inverse problems, [55]. With this background in mind, and to set the stage for formulating our learning problem as learning an operator between two Hilbert Spaces, let us introduce a (regular)‘link function’ $\Phi:\mathbb{R}\to[m,\infty)$ such that any conductivity $\gamma\in\Gamma$ is given by $\gamma=\Phi\circ f$ for some $f\in L^{\infty}(\Omega)\cap L^{2}{(\Omega)}:=\tilde{\Gamma}$. An example of such a link function is given in [57, Example 8]. Now we will modify our operator learning problem. For this we first define a map, $\mathcal{G}^{\dagger}:D_{\Lambda}\to\tilde{\Gamma}$ where $\mathcal{G}^{\dagger}=\Phi^{-1}\circ G^{\text{true}}$. Thus knowing $\mathcal{G}^{\dagger}$ completely determines $G^{\text{true}}$. Next we extend this map: Let $\mathcal{G}^{\dagger}:D_{\Lambda}\to L^{\infty}(\Omega)\cap L^{2}(\Omega)$ be the map as above, then there exists an extension $\mathcal{G}^{\dagger}_{\text{ext}}:\mathbb{H}_{r}\to L^{\infty}(\Omega)\cap L^{2}(\Omega)$ such that $\mathcal{G}^{\dagger}_{\text{ext}}|_{D_{\Lambda}}=\mathcal{G}^{\dagger}$. The extension $\mathcal{G}^{\dagger}_{\text{ext}}$ exists by Lemma 7.4. The map $\mathcal{G}^{\dagger}_{\text{ext}}$ will be approximated by a deep neural operator called DeepONet. Essentially, a neural operator is a parametric map between the input space (e.g. space of N-t-D operators) and the output space (e.g. space of admissible conductivities) that can be interpreted as a composition of three maps, an encoder, an approximator and a reconstructor, see [44] and Fig. 1. As such, an upper bound on the error in approximating the true operator by a neural operator can be split into upper bounds on the encoder error, the approximator error and the reconstructor error correspondingly. Now we recall the components of a DeepONet as proposed in [44, 51] and adapt the definitions as applicable to the case of EIT. To that end, we define the following operators while keeping the terminology the same as in [44]:
[…]
Figure 1: The true map $\mathcal{G}^{\dagger}_{\text{ext}}$ is approximated by a composition of three maps, encoder $\mathcal{E}$, approximator $\mathcal{A}$ and reconstructor $\mathcal{R}$. The resultant error in the approximation thus comprises of encoder, approximator, and reconstructor errors.
  • Encoder: Given a set of $M$ electrodes covering the boundary of the domain, $E_{m}\subset\partial\Omega$, $m=\{1,\cdots,M\}$ we define the encoder:
    (8) $\mathcal{E}_{M}:=\mathcal{E}:D_{\Lambda}\subset\mathbb{H}_{r}\to\mathbb{R}^{M}\times\mathbb{R}^{M};\quad\tilde{\Lambda}_{\gamma}\mapsto\tilde{R}_{M}(\gamma)$
    The encoder can be thought of as ‘sensor evaluations’ in the spirit of [44]. Whenever it is clear from context, we will denote $\mathcal{E}_{M}$ by $\mathcal{E}$. It follows from Lemmas 7.5 and 7.6 that the map $\mathcal{E}$ is continuous and from Lemma 7.4 can be extended continuously on $\mathbb{H}_{r}$. With a slight abuse of notation, we call this extended map $\mathcal{E}$ too.
  • Decoder: While the encoder is a projection of an infinite-dimensional object (i.e. $\tilde{\Lambda}_{\gamma}\in\mathbb{H}_{r}$) into a finite-dimensional space, a decoder lifts back a finite dimensional representation, $R_{M}(\gamma)$ into its corresponding map in the space of Hilbert-Schmidt operators. This lifting is done using the construction given in [21, section 6] and results in an operator ${\Lambda}^{M}_{\gamma}$, see [21, eqn. 6.8]. Thus, we have:
    (9) $\mathcal{D}_{M}:=\mathcal{D}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{H}_{r};\quad\tilde{R}_{M}(\gamma)\mapsto\tilde{\Lambda}^{M}_{\gamma}.$
    Similar to above, whenever it is clear from context, we will use the shorthand notation $\mathcal{D}$ instead of $\mathcal{D}_{M}$. Note that, $\mathcal{D}\circ\mathcal{E}(\tilde{\Lambda}_{\gamma})=\tilde{\Lambda}^{M}_{\gamma}.$ In particular, we use here as the decoder the construction outlined in [5, Section 3]. Indeed, given any Neumann data $f\in L^{2}_{\diamond}(\partial\Omega)$, $\tilde{\Lambda}^{M}_{\gamma}f:=\tilde{R}_{M}\tilde{P}^{E}_{M}{f}$ where the map $\tilde{P}^{E}_{M}$ is a (auxiliary) projection operator defined in accordance with [5, eq. (5.1)].
  • Approximator: The approximator is a deep (feedforward) neural network defined as:
    (10) $\mathcal{A}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}^{P};\quad\tilde{R}_{M}(\gamma)\mapsto\{\mathcal{A}_{k}\}_{k=1}^{P}$
    The composition of encoder and approximator, i.e. $\beta:=\mathcal{A}\circ\mathcal{E}:\mathbb{H}^{r}\to\mathbb{R}^{P}$ will be referred to as the branch net.
  • Reconstructor: Following [44], we denote a trunk-net:
    (11) $\mathbf{\tau}:\mathbb{R}^{d}\to\mathbb{R}^{P+1},\quad y=(y_{1},\cdots,y_{d})\mapsto\{\tau_{k}\}_{k=0}^{P}$
    where each $\tau_{k}$ is a feed-forward neural net. This can be thought of as a point-encoder which encodes a point $y\in\mathbb{R}^{d}$ into a $P+1$ dimensional representation. Subsequently we denote a $\mathbf{\tau}$-induced reconstructor:
    (12) $\mathcal{R}_{P}:=\mathcal{R}:\mathbb{R}^{P}\to L^{2}(\Omega),\quad\mathcal{R}(\alpha_{k}):=\tau_{0}(y)+\Sigma_{k=1}^{P}\alpha_{k}\tau_{k}(y).$
  • Projector: Given a reconstructor $\mathcal{R}$ as above with $\tau_{k}\in L^{2}(\Omega);k=0,\cdots,P$ and $\{\tau_{k}\}_{k=1}^{P}$ as linearly independent we can define a projector:
    (13) $\mathcal{P}_{P}:=\mathcal{P}:L^{2}(\Omega)\to\mathbb{R}^{p},\quad\mathcal{P}(f)=\left(\langle f-\tau_{0},\tau_{1}^{*}\rangle,\cdots,\langle f-\tau_{0},\tau_{P}^{*}\rangle\right)$
    where $\tau_{1}^{*},\cdots,\tau_{P}^{*}$ denotes the dual basis of $\tau_{1},\cdots,\tau_{p}$ such that:
    $\langle\tau_{k},\tau_{l}^{*}\rangle=\delta_{kl}\quad\forall\text{}k,l\in\{1,\cdots,P\}.$
Figure 2: DeepONet architecture used in this manuscript. It feeds the EIT measurements into a Branch Net and 2d point on $\Omega$ into a Trunk Net. The outputs of both networks are combined via an inner product to predict the $\sigma_{x,y}$ at points $(x,y)$.
3 Main Approximation Theorem
The quality of the approximation will be determined by the extent to which the maps in Fig. 1 commute. The neural operator $\mathcal{G}_{\theta}$ will approximate the ‘true map’ $\mathcal{G}^{\dagger}_{\text{ext}}$. Recall that we denote encoding (by sensor evaluations) by $\mathcal{E}$, decoding by $\mathcal{D}$, Reconstructor map by $\mathcal{R}$, the projector map by $\mathcal{P}$, and the approximator map between finite dimensional spaces as $\mathcal{A}$. Referring to Fig. 1, notice that our neural operator will have the form $\mathcal{G}_{\theta}=\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}$. Our goal is to find an upper bound on the error incurred in approximating $\mathcal{G}^{\dagger}_{\text{ext}}$ by $\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}:=\mathcal{G}_{\theta}$ over the space of shifted N-t-D operators corresponding to admissible conductivities. More particularly we will prove the following theorem.
Theorem 3.1.
Consider the sets $\mathbb{H}_{r},D_{\Lambda}$, $\tilde{\Gamma}$, and $\mathcal{G}^{\dagger}_{\text{ext}}$ as above. Note $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$. We will show that $\forall\epsilon>0$, there exist numbers $M\text{ and }P$ and continuous maps $\mathcal{E}:\mathbb{H}_{r}\to\mathbb{R}^{M}\times\mathbb{R}^{M}$, $\mathcal{R}:\mathbb{R}^{P}\to\tilde{\Gamma}$ and $\mathcal{A}:\mathbb{R}^{M}\times\mathbb{R}^{M}\to\mathbb{R}^{P}$ such that:
(14) $\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert\mathcal{G}^{\dagger}_{\text{ext}}(\tilde{\Lambda}_{\gamma})-\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}(\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}\leq\epsilon$
Proof 3.2.
The error term can be split in the following way,
$\sup_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert$ $(\mathcal{G}_{\theta}-\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}=\sup_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert(\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}-\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}$
$=\sup_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert(\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}+\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}-\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}$
$\leq\sup_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\bigg{(}\underbrace{\lvert\lvert(\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}}_{E_{\mathcal{A}}}$
$+\underbrace{\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}}_{E_{\mathcal{E}}}$
(15) $+\underbrace{\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}-\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}}_{E_{\mathcal{R}}})\bigg{)}$
where $E_{\mathcal{A}},E_{\mathcal{E}}$ and $E_{\mathcal{R}}$ denote the approximator error, encoder error, and the reconstructor error respectively. In the next three subsections we will analyze each of these error terms separately. In particular, we will show that for any given $\epsilon$ the network parameters can be chosen to ensure that each of the error terms separately is of the order $\epsilon/3$. Here and below, the symbol $\lesssim$ means that the inequality holds upto some constant factor, and the notation $\lVert\cdot\rVert_{L^{2}}$ is a shorthand for $\lVert\cdot\rVert_{L^{2}(\Omega)}$. Furthermore, the fact that $\Omega$ is a smooth bounded domain allows us to translate upper bounds in $L^{\infty}(\Omega)$ easily into upper bounds with respect to $L^{2}(\Omega)$ norm, we will use this fact freely wherever needed.
3.1 Approximator error
First we estimate the approximation error, $E_{\mathcal{A}}$. To that end, we examine the continuity of various maps appearing in the error terms above. Recall that $\tilde{\Lambda}_{\gamma}$ is a bounded map for admissible conductivities, [27, Theorem A.3]. As the map $\mathcal{R}$ is similar to the one defined in [44], the Lipschitz constant of $\mathcal{R}$ follows from [44, Lemma 3.2]. Consider the encoding map $\mathcal{E}(\tilde{\Lambda}_{\gamma}):=\tilde{R}^{M}_{\gamma}$, then we can show that this is Lipschitz, too. This follows from the discussion in [45, Sections 2 and 4] from which we first get that the map $\gamma\mapsto\tilde{R}^{M}_{\gamma}$ is Lipschitz and subsequently we apply the estimate ([28, Theorem 2.3]), $\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}\leq C\lvert\lvert\tilde{\Lambda}_{\gamma_{1}}-\tilde{\Lambda}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}$. As $D_{\Lambda}$ is a compact set from Lemma 7.1, then $\mathcal{E}(D_{\Lambda})$ is also compact. Now the map, $\mathcal{G}^{\dagger}_{\text{ext}}$ is a Lipschitz continuous map for piecewise analytic conductivities, [28, Theorem 2.3]. Furthermore, the map $\mathcal{P}$ is Lipschitz by similar argument as in [44, Lemma 3.2, 3.3]. $\mathcal{D}$ is also Lipschitz from the following estimate and the fact that the auxiliary projection operator, $\tilde{P}^{E}_{M}$ is bounded, see e.g. [5].
$\lvert\lvert(\tilde{\Lambda}^{M}_{\gamma_{1}}-\tilde{\Lambda}^{M}_{\gamma_{2}})f\rvert\rvert_{L^{2}_{\diamond}(\partial\Omega)}\leq C(z,\phi,\Omega,d,M)\lVert\tilde{R}^{M}_{\gamma_{1}}-\tilde{R}^{M}_{\gamma_{2}}\rVert_{l^{2}}\lVert\tilde{P}^{E}_{M}f\rVert_{H^{s}(\partial\Omega)}.$
Thus the composition $\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}:=G:\mathbb{R}^{M\times M}\to\mathbb{R}^{P}$ is also Lipschitz. Here and below, let $\mathrm{Lip(\mathcal{R})}$ and $\mathrm{Lip(\mathcal{E})}$ denote the Lipschitz constants for the (Lipschitz) maps $\mathcal{R}$ and $\mathcal{E}$. Thus
$E_{\mathcal{A}}$ $\leq\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert(\mathcal{R}\circ\mathcal{A}\circ\mathcal{E}-\mathcal{R}\circ G\circ\mathcal{E})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}$
$\leq\operatorname{Lip}(\mathcal{R})\cdot\sup_{X\in\mathcal{E}(D_{\Lambda})}\left\|\mathcal{A}(X)-G(X)\right\|,$
The upper bound on the approximation error can be evaluated by finding an upper bound for the error incurred in approximating the map $G:\mathbb{R}^{M\times M}\to\mathbb{R}^{P};{}X\mapsto(G_{1}(X),\cdots,G_{P}(X))$ by a neural network $\mathcal{A}$. This can be done using the results proved by Yarotsky in [75]. To make this precise, we begin by approximating each of the maps $G_{j}:\mathbb{R}^{M\times M}\to\mathbb{R},{}j\in\{1,\cdots,P\}$ by a neural network $\mathcal{A}_{j}:\mathbb{R}^{M\times M}\to\mathbb{R}$ and then combining all the individual approximations into a single neural network $\mathcal{A}$ wherein
$\mathrm{size}(\mathcal{A})\leq P\max_{j=\{1,\cdots,P\}}\mathrm{size}(\mathcal{A}_{j}).$
Then it follows from [75, Theorem 1] (see also discussion following [44, Theorem 3.17])that for any given $\epsilon/3:=\tilde{\epsilon}>0$, there exists a neural network $\mathcal{A}_{j}$ with
$\mathrm{depth}(\mathcal{A}_{j})\lesssim(1+\log(\epsilon^{-1}))\quad\mathrm{size}(\mathcal{A}_{j})\lesssim\tilde{\epsilon}^{-m}(1+\log(\tilde{\epsilon}^{-1}))$
such that
(16) $\sup\limits_{X\in\mathcal{E}(D_{\Lambda})}\lVert G_{j}(X)-\mathcal{A}_{j}(X)\rVert\leq\tilde{\epsilon}.$
Thus there exists a neural network $\mathcal{A}$ whose size is of the order $\mathcal{O}(P\tilde{\epsilon}^{-M})$ that can approximate a Lipschitz continuous mapping $G:\mathbb{R}^{M\times M}\to\mathbb{R}^{P}$ within a given $\epsilon/3=\tilde{\epsilon}$ error, see also [44, section 3.6.1].
3.2 Encoding error
For shifted N-t-D operators $\tilde{\Lambda}_{\gamma}$ corresponding to admissible conductivities, the encoding error is given by:
(17) $\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}=\lvert\lvert\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}((\mathcal{D}\circ\mathcal{E}-\mathbb{I}_{\mathbb{H}_{r}})\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}$
where $\mathbb{I}_{\mathbb{H}_{r}}$ is the identity map on $\mathbb{H}_{r}$. Observe that $(\mathcal{D}\circ\mathcal{E}-\mathbb{I}_{\mathbb{H}_{r}})\tilde{\Lambda}_{\gamma}=\tilde{\Lambda}^{M}_{\gamma}-\tilde{\Lambda}_{\gamma}$. From [21, Corollory 6.2], we know that by suitably shrinking the size of electrodes as their number $M$ grows we get
Lemma 3.3.
[5, Theorem 7.8] For $M\in\mathbb{N}$ denoting the number of electrodes,
(18) $\lvert\lvert\tilde{\Lambda}_{\gamma}-\tilde{\Lambda}^{M}_{\gamma}\rvert\rvert_{\mathcal{L}(L^{2}_{\diamond}\to L^{2}_{\diamond})}:=\delta_{M}\to 0\quad\text{as}\quad M\to\infty.$
As a simple consequence to the lemma above, and similar to the calculations in [1, Lemma 5] one can establish the relation between operator norm $\mathcal{L}(L^{2}_{\diamond}\to L^{2}_{\diamond})$ and the information-theoretically relevant $\mathbb{H}_{r}$-norm to get $\lvert\lvert\tilde{\Lambda}_{\gamma}-\tilde{\Lambda}^{M}_{\gamma}\rvert\rvert_{\mathbb{H}_{r}}\to 0$ as $M\to\infty$. As is shown in Lemma 7.1 below, $D_{\Lambda}$ is a compact set of $\mathbb{H}_{r}$. Consider the finite rank projection operators, $U_{M}:\mathbb{H}_{r}\to\mathbb{H}_{r}$ where $U_{M}:=\mathcal{D}_{M}\circ\mathcal{E}_{M}$ where $\mathcal{E}_{M}$ and $\mathcal{D}_{M}$ are as defined in section 2. From Lemma 3.3, we have that for any $\tilde{\Lambda}_{\Gamma}\in D_{\Lambda}$:
(19) $\lim\limits_{M\to\infty}\sup\limits_{\tilde{\Lambda}_{\gamma}}\lvert\lvert\Lambda_{\gamma}-U_{M}(\tilde{\Lambda}_{\gamma})\rvert\rvert_{\mathbb{H}_{r}}=0.$
Let, $Z:=\bigcup\limits_{M\in\mathbb{N}}U_{M}(\tilde{\Lambda}_{\gamma})\cup D_{\Lambda}$. Then $Z$ is a compact set of $\mathbb{H}_{r}$ from [42, Lemma 46]. Thus $\mathcal{G}^{\dagger}_{\text{ext}}:Z\to\tilde{\Gamma}$ is uniformly continuous on $Z$ and there exists a modulus of continuity, $\omega$ such that:
(20) $\forall z_{1},z_{2}\in Z,\quad\lvert\lvert\mathcal{G}^{\dagger}_{\text{ext}}(z_{1})-\mathcal{G}^{\dagger}_{\text{ext}}(z_{2})\rvert\rvert_{L^{\infty}}\leq\omega(\lvert\lvert z_{1}-z_{2}\rvert\rvert_{\mathbb{H}_{r}})$
On the other hand, given any $\epsilon$, we can clearly choose an $M:=M(\epsilon)$ such that
(21) $\sup\limits_{z\in D_{\Lambda}}\omega(\lvert\lvert U_{M}(z)-z\rvert\rvert_{\mathbb{H}_{r}})\leq\epsilon/3.$
Finally, we know from arguments above that the maps $\mathcal{R}$ and $\mathcal{P}$ are Lipschitz continuous, see subsection 3.1. This shows that by using a large enough number of electrodes $M$, the relative CEM measurements can be used to approximate the continuum relative N-t-D operator to any arbitrary precision. More particularly, combining eqns.(20) (21) and the fact that $\mathcal{R}$ and $\mathcal{P}$ are Lipschitz we get that given any $\epsilon>0$, there exists $M:=M(\epsilon)$ such that,
(22) $\sup\limits_{\tilde{\Lambda}_{\gamma}\in D_{\Lambda}}\lvert\lvert(\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}}\circ\mathcal{D}\circ\mathcal{E}-\mathcal{R}\circ\mathcal{P}\circ\mathcal{G}^{\dagger}_{\text{ext}})\tilde{\Lambda}_{\gamma}\rvert\rvert_{L^{2}}\lesssim\epsilon/3.$
3.3 Reconstruction error
Note that the output of the DeepONet lies in the space of $L^{2}$ functions. For any $\tilde{\Lambda}_{\gamma}\in D_{\Lambda}$, for the third term in (3.2) we have, $E_{\mathcal{R}}\leq\sup\limits_{\tilde{\Lambda}\in D_{\Lambda}}{\lvert\lvert(\mathcal{R}\circ\mathcal{P}-\mathbb{I}_{\Omega})(\mathcal{G}^{\dagger}_{\text{ext}}\tilde{\Lambda}_{\gamma})\rvert\rvert_{L^{2}}}$, where $\mathbb{I}_{\Omega}$ represents the identity map on the space of $L^{2}(\Omega)$ functions. Let $\tilde{\Lambda}_{\gamma}\in D_{\Lambda}$ and denote $g:=\mathcal{G}^{\dagger}_{\text{ext}}(\tilde{\Lambda}_{\gamma})=\mathcal{G}^{\dagger}(\tilde{\Lambda}_{\gamma})=\Phi^{-1}\circ\gamma$, where $\Phi$ is the link function as described in section 2. Since $\gamma$ is a piecewise analytic function, hence it belongs to some $H^{s}(\Omega)$ for $20<s<1/$ 2. It is easy to see that,
$\mathcal{R}\circ\mathcal{P}(g)=\tau_{0}^{\perp}+\Sigma_{k=1}^{P}\langle g,\tau_{k}\rangle\tau_{k}$
where $\tau_{0}^{\perp}:=\tau_{0}-\Sigma_{k=1}^{P}\langle\tau_{0},\tau_{k}^{*}\rangle\tau_{k}$ is the projection of $\tau_{0}$ onto the orthogonal complement of $\mathrm{span}(\tau_{1},\cdots,\tau_{p})$. Following arguments from [44, Lemma 3.2,3.3, 3.4], we have the following Lemma.
Lemma 3.4.
For any $10<\epsilon<$ 1 and some $C_{1}$ such that $\lVert g\rVert_{H^{s}}\leq C_{1}$, there exists a trunk net $\tau:\mathbb{R}^{d}\to\mathbb{R}^{P}$ with
$\mathrm{size}(\mathbf{\tau})\sim 1+P\log(P/(\epsilon/3))^{2}$
$\mathrm{depth}(\mathbf{\tau})\sim 1+\log(P/(\epsilon/3))^{2}$
such that
$E_{\mathcal{R}}\lesssim C_{1}P^{-s/d}+\epsilon/3.$
Here we can choose $P$ to be large enough such that, $E_{\mathcal{R}}\lesssim\epsilon/3.$
Finally, the upper bounds shown on the three kinds of error, viz. approximator error, encoder error and reconstructor error, prove Theorem 3.1.
4 Network Architecture
In this section, we describe the network architecture used in our experiments which corresponds to the one depicted in Fig. 2. The input for the Branch network are the measurements of the EIT experiment. For the numerical simulations, we imagine using 16 different electrodes through which the current will be injected and the voltages measured. In practical settings, current is injected by using a pair of adjacent source and sink electrodes. For each such current pattern, voltage differences are measured across adjacent electrodes on the remaining $14$ electrodes. Hence, we obtain $13$ voltage difference measurements per injected current pattern. In total, we get $16*13=208$ measurements which can be arranged in the form of a matrix. Many a times, one uses zero-padding to fill in the missing entries of the matrix to make it a square matrix. The measurements are then fed into a series of fully connected ResNet blocks to output a vector of the same dimensionality as the output of the later defined Trunk network (See Table LABEL:BranchNet).
Table 1: Branch Network Architecture
Network Block Input Size Output Size
ResNet 208 1024
ResNet 1024 1024
ResNet 1024 1024
ResNet 1024 512
ResNet 512 512
ResNet 512 512
ResNet 512 512
The Trunk network takes points $(x,y)\in\mathbb{R}^{2}\subset\Omega$ as input. It then uses a series of ResNet blocks to get an output of the same dimensionality as the Branch network (See Table LABEL:TrunkNet). An inner product operator is used to combine both outputs into a prediction $\hat{\sigma}_{x,y}\in\mathbb{R}$. $\hat{\sigma}_{x,y}$ is to be interpreted as a prediction of $\sigma$ at point $(x,y)\in\mathbb{R}^{2}\subset\Omega$ given the measurements.
Table 2: Trunk Network Architecture
Network Block Input Size Output Size
ResNet 2 256
ResNet 256 512
One interesting observation from the computational perspective is that, for each EIT experiment, the Branch network only has to be computed once. The Trunk network has to be evaluated for many points (in parallel). That is why the Trunk network is chosen to be much more lightweight than the Branch network. We note that we experimented and evaluated our method with many different network architectures of similar sizes and they all performed reasonably well. We chose to present this setup as this seemed to do the best in our experiments.
5 Experiments
In this section, we show how our proposed Brunch-Trunk network performs in computational experiments.
5.1 Training the Neural Network
5.1.1 Training Data
We used a finite element based PDE solver to generate a dataset of $16,384$ data pairs consisting of 1, 2, 3 and 4 anomalies and the corresponding measurements. That is we used $4096$ conductivity’s and the corresponding measurements for each 1, 2, 3, and 4 circular anomalies. The measurements where obtained with a finite element based PDE solver[50] solving the equation 1.
During training we generate a point cloud 1024 points $(x,y)\in\mathbb{R}^{2}\subset\Omega$ and the corresponding $\sigma_{x,y}$. Our sampling is to be understood as sampling $50\%$ of the points from the background and $50\%$ of the points from the anomalous regions. This is only done if there are anomalies in $\sigma$. The reason for this kind of sampling is that in our experience it significantly increased the correct detection on small anomalies. The idea behind this is that if only a very small amount of samples is anomalous the neural network has the tendency of predicting everything as background. We note that more sophisticated sampling methods have been advocated in the literature. However, for the verification of our theoretical results, this simple sampling scheme tends to also work well.
5.1.2 Training and Testing
Our neural network is of the form shown in Figure 2, consisting of the two components defined in Table LABEL:BranchNet and LABEL:TrunkNet. During the Training process, we first evaluate the measurement data via our Branch network followed by evaluating the corresponding large set of 1024 $2D$ points with the Trunk network. Then the inner product of both outputs is computed. Each estimated value $\hat{\sigma}_{x,y}$ is then compared with the ground truth $\sigma_{x,y}$ with a weighted- $L^{2}$-loss. The weighted $L^{2}$ loss is used to reflect the fact that during training we trained our network using $50\%$ points from the anomalous area and the rest from the background area. Furthermore, during the training process, we use a Batch size of 32 conductivity setups. The training is performed for 200 epochs using the ADAM optimizer. The 1-D cut plots in Figure 3 denotes the values of conductivity function along a 1-D profile indicated by a white line in the original phantom.
5.2 Performance
In this subsection, we describe the performance of our proposed method in our numerical studies and compare it against a standard method used for EIT inversion.
5.2.1 Baseline: Iteratively Regularized Gauss-Newton Method
We use the Iteratively Regularized Gauss-Newton (IRGN) method, see e.g. [3], as a baseline for our approach. The IRGN method is a popular gradient-based optimization method for solving the EIT problem in an analytical setting on a PDE mesh. We use the standard Tikhonov regularized IRGN method, see [3] for a good description of this method. We do not provide an extensive comparison with other methods because our main results are on the theoretical side. Hence, the goal is to show that our method is competitive with one state-of-the-art method.
5.2.2 Comparison
We evaluate both methods on 5 images at $1\%$ additive measurement noise. The comparison results can be found in Figure 3. On the left-hand side is the ground truth image. The second row is the IRGN method followed by our proposed method. Note the first three figures are in the same color range. The rightmost image is a $1D$ cut through the previous 3 images. The cut direction is visualized in a white line in the ground truth image. We observe that both methods are approximately finding the correct location of the anomaly. However, our method outperforms the IRGN method in terms of obtaining the correct sigma in the anomalous areas. Some of the numerical values can be found in noise study in Table LABEL:noise_table.We also note the good generalization ability of the proposed network in practice. In particular, observe that in the third and fourth row of Fig 3, we attempt to reconstruct thin and stretched rectangular shaped anomalies, while the network was trained only on circular anomalies.
Ground Truth           IRGN Method           DeepONet               1-D Cut
Figure 3: Comparison of reconstructions of our DeepONet with the classical IRGN method. We observe that the DeepONet obtains much closer $\sigma$ values to the ground Truth in the anomalous regions. The white line in the ground truth is the cut at which the plot on the right hand side is displayed.
5.3 Weighted $L^{2}$ error
We believe that the standard $L^{2}$ error is not the correct metric because, for images with small anomalous regions, a small error does not imply that the location and conductivity value of the anomaly were estimated properly. For example, in such a case, the reconstruction where everything is exactly the background value, would have relatively low $L^{2}$ errors. However, the reconstruction would not have any practical use. Hence, we use the weighted $L^{2}$ error that takes $50\%$ weight of the ground truth anomalous areas and the rest from the background area. In this metric, the DeepONet method outperforms the IRGN method as expected from seeing the images as well as the error table below.
5.4 Noise Study
The stability of the method against higher noise levels on a single image is evaluated. All experiments of the errors are repeated ten times (On Fig. 3, Row 1) and the mean and standard deviation are displayed in Table LABEL:noise_table. Overall, the weighted $L^{2}$-Error is increasing in a noise pattern in the DeepONet method. The same is true for the IRGN method. However, the IRGN method does have lower errors. We note that for large noise levels from $7\%$, the reconstructed anomalies find significantly worse locations and shapes.
Table 3: Noise Study - Weighted $L^{2}$ Fig. 3, Row 1
Noise Level $1\%$ $2\%$ $3\%$ $5\%$ $9\%$
DeepONet $0.54\pm 0.52$ $0.63\pm 0.66$ $0.65\pm 0.70$ $1.05\pm 1.25$ $1.62\pm 2.17$
IRGN $14.16\pm 0.79$ $14.15\pm 0.98$ $14.34\pm 2.87$ $14.65\pm 3.75$ $14.21\pm 4.71$
6 Conclusion
In this work, we have generalized the earlier known approximation-theoretic results for approximating a map between infinite-dimensional function spaces using DeepONets to the case where we can provide similar guarantees even when the goal is to learn a map between a space of (N-t-D) operators to the space of conductivities (functions). This map arises naturally in the framework of Electrical Impedance Tomography which is a non-invasive medical imaging modality that is becoming increasingly prominent in especially several neuro-imaging applications. As deep-learning networks are becoming increasingly popular, it behooves us to go beyond black-box description of such networks and to make them interpretable so as to facilitate deployment with confidence in critical applications such as medical imaging. Providing approximation-theoretic guarantees of the kind given in this work is a crucial first step in achieving the goal of interpretable neural-networks. In fact, we believe that the analysis carried out in this work can be easily adapted to many other problems wherein the input and output spaces are classes of operators satisfying certain general assumptions. We have compared our approach against the widely used IRGN benchmark and show that the proposed network outperforms the baseline in numerical studies. We also performed a noise study and show that robustness of the deep-learning based against a weighted $L^{2}$ error metric as introduced in this work. To our knowledge, this is the first work that uses a DeepONet architecture for approximating an operator-to-function map as appearing in EIT, while also providing a theoretical guarantee to show its effectiveness. As already stated, our numerical results back-up our theoretical findings.
7 Appendix
In this section, we collect some auxiliary results needed for the proof of our main theorem.
Lemma 7.1.
The set $D_{\Lambda}$ is a compact set in $\mathbb{H}_{r}$.
Proof 7.2.
First of all we note that $\mathbb{H}_{r}$ is a separable Hilbert Space with a countable orthonormal basis given in [1, section 5.1]. We recall the following theorem which is given in e.g. [15, Chapter 2], [65, Theorem 197].
Theorem 7.3.
[65, Theorem 197] Let $H$ be a separable Hilbert space, and let $\{e_{k}\}_{k=1}^{\infty}$ be an orthonormal basis of $H$. Then a subset $K\subset H$ is compact iff:
  1. 1.
    $K$ is closed,
  2. 2.
    $K$ is bounded, and
  3. 3.
    $K$ has equi-small tails with respect to $\{e_{k}\}$, i.e., for every $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for all $\vec{v}\in K$,
    $\sum_{k=N}^{\infty}|\langle\vec{v},e_{k}\rangle|^{2}<\varepsilon^{2}.$
Now, the fact that $D_{\Lambda}$ is bounded follows from [21, (2.5)] as before. The fact that $D_{\Lambda}$ has equi-small tails follows from [1, Lemma 4] in which we choose $J,K$ and $\nu$ such that $C\min(J,K)^{-\nu}\leq\varepsilon$ for any given $\varepsilon$. Thus it remains to show that $D_{\Lambda}$ is closed. To that end, choose a sequence of relative N-t-D operators $(\tilde{\Lambda}_{\gamma_{n}}:\gamma_{n}\in\Gamma)$ such that they converge to some operator $T$ in $\mathbb{H}_{r}$ i.e., $\tilde{\Lambda}_{\gamma_{n}}\to T$. We will show that there exists some $\gamma\in\Gamma$ such that $T=\tilde{\Lambda}_{\gamma}$ i.e. $T$ is also a N-t-D operator. First of all note that since $(\tilde{\Lambda}_{\gamma_{n}})$ converges, hence $(\tilde{\Lambda}_{\gamma_{n}})$ is actually Cauchy in the $\mathbb{H}_{r}$ norm. Using the relation between the operator norm and the information-theoretic Hilbert-Schmidt norm as stated in [1, Lemma 5], we can show that $(\Lambda_{\gamma_{n}})$ is also Cauchy in the operator norm $\mathcal{L}(L^{2}\to L^{2})$. By using the inverse continuity result for the Calderon problem, see [28, Theorem 2.3], we get that the sequence $\gamma_{n}$ is Cauchy with respect to the $L^{\infty}$ norm and hence converges, say $\gamma_{n}\to\gamma_{0}\in L^{\infty}(\Omega)\cap\Gamma$. Let us now consider, $\Lambda_{\gamma_{0}}$. From forward continuity result for the Calderon problem, see [28, Lemma 2.5], $\Lambda_{\gamma_{n}}\to\Lambda_{\gamma_{0}}$. Now from uniqueness of the limit, $\Lambda_{\gamma_{0}}=T$. Hence the set, $D_{\Lambda}$ satisfies all three criteria for it be a compact subset of $\mathbb{H}_{r}$.
Lemma 7.4.
[17, Theorem 4.1] Let $X$ be an arbitrary metric space, $A$ a closed subset of $X$, $L$ a locally convex linear space, and $f:A\to L$ be a continuous map. Then there exists an extension $F:X\to L$ (which is continuous by definition) such that $F(a)=f(a)$ for all $a\in A$.
Finally, consider the encoding map $\mathcal{E}(\tilde{\Lambda}):=\mathcal{E}_{M}(\tilde{\Lambda})=\tilde{R}^{M}_{\gamma}.$ We have the following continuity results:
Lemma 7.5.
[45, section 2] The map $T:\gamma\to\tilde{R}_{\gamma}^{M}$ is Lipschitz continuous, i.e. $\lvert\lvert\tilde{R}^{M}_{\gamma_{1}}-\tilde{R}^{M}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}\leq\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}$
Lemma 7.6.
[28, Theorem 2.3] $\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{2}}\leq c\lvert\lvert\gamma_{1}-\gamma_{2}\rvert\rvert_{L^{\infty}}\leq C\lvert\lvert\tilde{\Lambda}_{\gamma_{1}}-\tilde{\Lambda}_{\gamma_{2}}\rvert\rvert_{\mathcal{L}(L^{2}\to L^{2})}$
From previous two Lemmas, we get that the map $\mathcal{E}$ is Lipschitz continuous.
References Cited